OpenStudy (anonymous):

Please I'll Give Medal!! (7x^3 +2x^2 + (6x^3 + 2x^2)

OpenStudy (anonymous):

\[(7x^3 +2x^2 + 6x^3 + 2x^2) = (7x^3 + 6x^3 +2x^2 + 2x^2)\] \[ (7x^3 + 6x^3 +2x^2 + 2x^2) = (7 + 6) \times x^3 + (2 + 2) \times x^2\] Can you do the rext yourself?